OpenStudy (anonymous):

What polynomial has roots of –6, –4, and 1 ?

hartnn (hartnn):

(x+6)(x+4)(x-1) simplify it by opening brackets....

OpenStudy (anonymous):

Can you show me

hartnn (hartnn):

use this (a+b)(c+d)=(ac+ad+bc+bd)

OpenStudy (anonymous):

Show me

hartnn (hartnn):

ok ok, \[(x+6)(x+4)(x-1)=(x^2+10x+24)(x-1)=x^3+9x^2+14x-24\] thats your final answer